# MIRAGE: KV Cache Optimization through Parameter Remapping for Multi-tenant LLM Serving

### v1

@@ 3.1. Limitations of CPU Offloading
3.2. Limitations of Swapping KV Cache
With emerging hardware offering higher CPU–GPU bandwidth, KV cache for non-executing (future) layers can be temporarily swapped into CPU memory and brought back to GPU memory when needed. This approach enables larger batch sizes and enhances overall system performance, as demonstrated by Pie (Xu et al., 2024). However, KV cache swapping presents significant challenges, primarily because the KV cache must be updated after every token generation (Yu et al., 2022; Kwon et al., 2023; Agrawal et al., 2024b). This results in frequent synchronization between CPU and GPU memory, introducing overhead that affects serving latency. In addition to synchronization overhead, the bidirectional nature of data transfers can reduce effective CPU–GPU bandwidth, preventing the system from achieving an ideal pipeline of data transfer and computation. Our measurements of CPU host memory (LPDDR5X [^5] The trend observed in our measurements also holds for other memory technologies (e.g., DDR4, DDR5) beyond LPDDR5X (Esmaili-Dokht et al., 2024).) bandwidth show that unidirectional (read-only of CPU host memory) CPU-to-GPU transfers can reach $\sim$427 GB/s. However, when the read-to-write ratio is 1:1, bandwidth drops to $\sim$366 GB/s, resulting in a $\sim$15reduction. This bandwidth degradation can further impact serving throughput.
3.3. Our Proposal: Parameter Remapping
We introduce Parameter Remapping, where, instead of swapping KV cache, we propose to dynamically remap the space used for model parameters to automatically and elastically expand the KV cache size at runtime. Dynamic parameter remapping reclaims GPU memory by repurposing a portion of the model parameters for KV cache usage during runtime, enabling online inference serving systems to handle peak loads when GPU memory becomes insufficient for KV cache. As shown in the two examples in Figure 4, deploying two OPT-13b models on an H100 GPU using either temporal (Figure 4) or spatial (Figure 4) sharing one GPU consumes approximately 52GB of the available 80GB memory for model parameters (Kwon et al., 2023), leaving only about 28GB for KV cache. By remapping one-third of the model parameters as KV Cache—either $2/3$ from the inactive model in temporal sharing or $1/3$ from each model in spatial sharing—the maximum batch size can be increased by up to 1.75×1.75$\times$1.75 ×, resulting in a corresponding 1.75×1.75$\times$1.75 × improvement in system throughput. Unlike KV cache swapping, parameter remapping involves non-blocking, unidirectional data transfers and does not require backing up model parameters from GPU to CPU, as parameters remain unchanged during inference. As a result, synchronization overhead is significantly reduced. Additionally, parameter remapping is particularly effective in multi-tenant LLM inference scenarios, such as multi-agent workflows (Talebirad and Nadiri, 2023; Chan et al., 2023; Li et al., 2024b; Han et al., 2024b; Kim et al., 2025) and production deployments where many models remain idle for long periods of time (Yu et al., 2025). In such settings, memory allocated to inactive models can be reclaimed for active ones, a flexibility that KV cache swapping lacks, as inactive models may have no KV cache to be swapped.
(a) CPU Compute.
(b) GPU Compute.
@@ 3.4. Challenges in Parameter Remapping
While parameter remapping can improve performance, it introduces potential overhead and requires careful design. To avoid impacting inference latency and throughput, parameter loading must ideally overlap with GPU computation, ensuring that the parameters of each layer are fully loaded before its execution begins. For example, while the GPU is executing layer 1, the parameters for layer 2 can be loaded concurrently from CPU to GPU memory. The ideal case is when layer 2 finishes loading before layer 1 finishes execution, allowing GPU memory to hold only two layers at a time (Sheng et al., 2023; Rasley et al., 2020), unlike approaches that preload the entire model into GPU memory. Although GH200 offers significantly higher CPU–GPU bandwidth than previous architectures, it remains insufficient for full on-demand remapping, as the execution of a layer is faster than transferring its parameters from CPU to GPU memory. As a result, only a fraction of the parameters—defined by the remapping percentage—can be dynamically transferred at runtime, while the rest must be preloaded into GPU memory before the inference serving starts. In this section, we examine the challenges involved in making these design decisions.
@@
Which models to remap? In multi-tenant LLM inference serving systems, not all models are active simultaneously (Patke et al., 2024; Yu et al., 2025), which presents an opportunity for the parameter remapping engine to prioritize remapping parameters of inactive models. However, this approach requires careful design, as overly aggressive offloading may lead to starvation—especially given the unpredictable nature of request arrival patterns in online serving environments (Stojkovic et al., 2024b). An inefficient design is likely to prevent certain models from responding promptly; for example, a suboptimal model selection can lead to a 22.0% increase in serving tail latency (§ 7.6).
@@
How many layers to remap? Determining the optimal remapping percentage is challenging due to the dynamic nature of request arrival rates in real-world online inference serving (Stojkovic et al., 2024a; Patel et al., 2024; Stojkovic et al., 2024b). The fluctuations in request arrival rates can affect the GPU computation time, which in turn affects the percentage of parameters that can be transferred without performance degradation. We illustrate the need for dynamic adjustment of the remapping percentage using an example of serving OPT-13b with varying batch sizes on the ShareGPT dataset (Team, 2025). Figure 5 shows how the GPU computation time for generating one token in the decode phase varies with batch size, while parameter load time—determined solely by the remapping percentage—remains constant. The intersection points of these curves represent configurations where parameter loading is fully overlapped with computation, maximizing efficiency. As shown, the optimal remapping percentage varies with changes in the runtime batch size, reflecting shifts in system load. Using a static remapping percentage may result in suboptimal resource use: underutilizing GPU memory at off-peak load and causing compute stalls at peak load. For example, a 30% remapping percentage tuned for average batch size in ShareGPT results in CPU–GPU bandwidth becoming a bottleneck during off-peak workloads. Therefore, the remapping percentage must be dynamically adapted at runtime.
@@
Which layers to target? Once the remapping percentage is set, the next design challenge is determining which specific layers to remap. This remains non-trivial, as MIRAGE must decide how to allocate limited GPU memory across all layers—determining which layers remain permanently in GPU memory and which are transferred on the fly. To achieve an efficient pipeline, a layer selection strategy is needed to guide the layer-sharing mechanism in MIRAGE. A suboptimal layer selection can lead to a reduction in serving throughput by 12.7% (§ 7.5).
4. MIRAGE Overview
@@
We introduce MIRAGE, a Dynamic Remapping Engine for multi-tenant LLM inference serving systems that elastically adjusts the KV cache size by remapping parameter memory as KV cache memory. Figure 6 shows the MIRAGE architecture, and depicts how MIRAGE can be integrated into existing LLM inference serving systems.
4.1. Architecture of MIRAGE
@@
MIRAGE is designed to have three key components, marked in green color in Figure 6.
@@
Async Transfer Engine: The Async Transfer Engine is responsible for loading parameters and coordinating synchronization between GPU computation and CPU-GPU data transfer. When the Remapping Controller initiates parameter remapping, a portion of GPU memory originally allocated for model parameters is repurposed for KV cache usage. If the reclaimed parameters are later required for computation, the Async Transfer Engine fetches them back from CPU memory to GPU memory. By leveraging the deterministic execution order of layers within each decode iteration, the Async Transfer Engine overlaps CPU–GPU data transfers with ongoing GPU computation.
Figure 6. Overview of MIRAGE.
@@ 4.2. Workflow
Figure 6 depicts the workflow of the interaction between various components to achieve the goal of dynamic parameter remapping in MIRAGE. The Scheduler [^6] MIRAGE is compatible with a wide range of popular scheduling policies (e.g., either spatial or temporal sharing of GPU resources for multi-tenant LLM inference serving). The scheduling policy design itself is beyond the scope of this paper. selects which models to serve, referred to as active models. Once active models are chosen, the Memory Allocator updates the parameter and KV cache memory usage accordingly. Using both model information and KV cache utilization data, the Remapping Controller determines whether parameter memory needs to be remapped to expand KV cache capacity. If remapping is required, the Remapping Controller updates memory usage records in the Memory Allocator and triggers the Async Transfer Engine to initiate data transfer. The GPU then executes the LLM inference tasks. With support from MIRAGE, a portion of parameter memory in GPUs can be dynamically reclaimed at runtime, allowing the system to accommodate larger batches or longer sequences.
@@ 5. Remapping Controller
MIRAGE enables dynamic, runtime management of GPU memory through its Remapping Controller. Figure 7 illustrates an example execution workflow of the Remapping Controller across five phases, each representing a snapshot of model activity within an LLM serving system. Depending on runtime conditions, the Remapping Controller decides either to initiate remapping, continue remapping additional layers from the active or inactive models, or halt remapping altogether (reclaimed KV cache memory restored for parameter storage). In this section, we describe how Remapping Controller answers the key questions, including when to remap (§ 5.1), which models to remap (§ 5.2), how many (§ 5.3), and which layers to remap (§ 5.4). Putting it all together, we then present an algorithm that can seamlessly integrate MIRAGE into existing LLM inference serving frameworks, enabling dynamic remapping of model parameter memory to automatically and elastically expand KV cache capacity at runtime (§ 5.5).
@@ 5.1. When to Remap?
The Remapping Controller initiates parameter remapping when the KV cache capacity is exhausted. While unidirectional parameter remapping incurs lower overhead than bidirectional KV cache swapping, MIRAGE activates it only when necessary. As such, accurately determining both the initiation and termination points of remapping is critical. Given the dynamic nature of real-world production workloads (Cortez et al., 2017; Hadary et al., 2020; Shahrad et al., 2020; Stojkovic et al., 2024a; Patel et al., 2024; Stojkovic et al., 2024b), the Remapping Controller enables remapping during peak request periods to accommodate larger batch sizes. During non-peak periods, when KV cache space is sufficient, remapping is disabled to avoid the overhead of reloading previously evicted parameters into GPU memory. The reclaimed KV cache memory is then restored for parameter usage.
@@ 5.2. Which Models to Remap?
GPU sharing background: Given the memory capacity of emerging GPUs, enabling GPU sharing is essential to fully utilize their processing capabilities. Prior works have examined both temporal (Bai et al., 2020; Gujarati et al., 2020; Xiao et al., 2020; Fu et al., 2024; Patke et al., 2024) (example in Figure 4) and spatial (Dhakal et al., 2020; Romero et al., 2021; Lee et al., 2024b; Strati et al., 2024; Choi et al., 2022; Xu et al., 2022; Han et al., 2024a; Wang et al., 2024a) (example in Figure 4) sharing strategies for efficient machine learning inference serving on GPUs. Temporal sharing allocates the entire GPU to one model at a time, allowing for larger batches per model. This approach is especially useful in multi-agent workflows—where the output of one model becomes the input to another (Talebirad and Nadiri, 2023; Chan et al., 2023; Li et al., 2024b; Han et al., 2024b; Kim et al., 2025)—and in cloud production environments, where models often experience idle periods with no incoming requests (Yu et al., 2025). In contrast, spatial sharing divides GPU resources among multiple models simultaneously, allowing for faster response times at the expense of smaller batch sizes per model (Lee et al., 2024b; Duan et al., 2024). Each GPU resource sharing strategy is tailored to specific use cases and can be selectively employed by multi-tenant LLM inference schedulers. The Remapping Controller in MIRAGE is designed to be compatible with any scheduling policy, although the design of the scheduler itself is beyond the scope of this study.
Model selection strategy: We now present the model selection strategy to determine which models to remap parameter layers from. The Remapping Controller prioritizes remapping parameters from inactive models with the lowest priority, as determined by the scheduling policy, rather than from active ones. Under temporal GPU resource sharing, models scheduled for future execution are considered inactive. In contrast, under spatial sharing, all models are treated as active, requiring each to remap parameters. The number of layers to remap and their selection are determined using the strategies described in § 5.3 and § 5.4. If the scheduler does not specify model priorities, the Remapping Controller defaults to a round-robin policy, treating all models equally. In this case, it applies a Most Recently Used (MRU) strategy, remapping parameters from the most recently activated model to defer parameter transfer costs as far into the future as possible. This strategy is especially effective when future request patterns are unpredictable or when model priorities are unspecified.
@@ 5.3. How Many Layers to Remap?
To hide parameter transfer latency by overlapping it with GPU computation, the transfer from CPU to GPU must finish before the computation completes; otherwise, it becomes a performance bottleneck. Since both model parameter sizes and CPU–GPU bandwidth are known ahead of runtime, the transfer time per layer can be profiled offline and treated as prior knowledge by the online inference system. As a result, the Remapping Controller only needs to monitor GPU computation time, denoted as $T_{Compute}$, at runtime to determine whether additional parameters can be remapped. For an active model, $T_{Compute}$ refers to the time required to generate the current token during the decode phase. For an inactive model, it represents the duration of the prefill phase, which is more compute-intensive and typically longer, as no tokens have been generated yet (Patel et al., 2024; Agrawal et al., 2024b). The longer $T_{Compute}$ in the prefill phase allows more layers in inactive models to be remapped compared to when the model is active. Let $T_{T}$ denote the time required to transfer a single layer of parameters. To prevent parameter loading from becoming a new bottleneck, the number of remapped layers $N$ must satisfy the constraint $T_{T}\cdot N\leq T_{Compute}$.
@@ 5.4. Which Layers to Target?
Layer selection strategy: Understanding the layer selection strategy requires recognizing the circular execution pattern of layers in LLM inference, as shown in Figure 7. The auto-regressive nature of LLM inference makes each of the subsequent tokens execute the same layers again. Thus, LLM inference can be viewed as a circular execution of layers, where the execution of the last layer for the current token is succeeded by the execution of the first layer for the next token. The Remapping Controller uses a uniform-interval layer selection strategy, in which evenly spaced layers are chosen to share a common GPU memory region allocated for their parameters. These layers alternate access to the shared region—at any given moment, some layer parameters reside in GPU memory, while others remain in CPU memory. When a layer not currently in GPU memory is needed, its parameters are loaded into the shared region, replacing those of the previously loaded layer.
@@
Example: To provide a more intuitive explanation, we use the example of Model-C with 8 layers (Phase 2 of Figure 7). The Remapping Controller employs a uniform-interval layer selection strategy. For instance, when remapping the parameters of a single layer, Layers 1 and 5 are selected to share a common GPU memory region (the other 6 layers remain in GPU memory permanently), alternating between CPU and GPU memory. The uniform-interval layer selection strategy is designed to balance serving latency across successive token generations. In contrast, ignoring the circular nature of LLM inference—by treating each token generation as an independent execution—may lead to suboptimal choices, such as selecting Layer 1 and Layer 8. Although this configuration allows sufficient time to load the parameters of Layer 8 after executing Layer 1, it does not provide enough time to reload the parameters of Layer 1 before it is needed again. This demonstrates the benefit of uniform-interval selection in circular execution, which ensures that each remapped layer can be restored to GPU memory in time for reuse.
@@
Theoretical proof: Consider an n-layer model in which the Remapping Controller remaps the parameter memory of $\alpha$ layers for KV cache. Consequently, all $n$ layers must be scheduled within the GPU memory footprint of only $n-\alpha$ layers. Suppose $m$ layers are transferred between CPU and GPU during each token generation. We note $L_{1},L_{2},\ldots,L_{m}$ as the $m$ layers (remapped layers in Figure 7) that are transferred from CPU to GPU. We define $T_{T}$ as the time taken to load parameters of a single layer from CPU to GPU memory, and $T_{c}$ as the time spent on computation for a single layer on GPU [^7] LLM architectures often adopt layers of uniform size, both in terms of parameters and compute requirements (Agrawal et al., 2024a; Liakopoulos et al., 2025), resulting in uniform parameter transfer and computation latency across layers. Therefore, $T_{C}$ and $T_{T}$ are equal for all (hidden) layers.. Our layer selection policy aims to satisfy two key objectives: (1) ensure that parameter transfer latency does not become a performance bottleneck; and (2) minimize CPU-GPU data transfer traffic.
@@
Consider that there are $k_{1}$ layers between $L_{1}$ and $L_{2}$, and $k_{m}$ layers between $L_{m}$ and $L_{1}$ (for the next token), we have:

### v2

@@ 3.4. Challenges in Parameter Remapping
While parameter remapping can improve performance, it introduces potential overhead and requires careful design. To avoid impacting inference latency and throughput, parameter loading must ideally overlap with GPU computation, ensuring that the parameters of each layer are fully loaded before its execution begins. For example, while the GPU executes layer 1, layer 2’s parameters can be concurrently loaded from CPU to GPU memory. The ideal case is when layer 2 finishes loading before layer 1 finishes execution, allowing GPU memory to hold only two layers at a time (Sheng et al., 2023; Rasley et al., 2020), unlike approaches that preload the entire model into GPU memory. Despite GH200’s higher CPU-GPU bandwidth, it remains insufficient for full on-demand remapping, since layer execution outpaces parameter transfer. As a result, only a fraction of the parameters, defined by the remapping percentage, can be dynamically transferred at runtime, while the rest must be preloaded into GPU memory before the inference serving starts. In this section, we examine the challenges involved in making these design decisions.
@@
Which models to remap? In multi-tenant LLM inference serving systems, not all models are active simultaneously (Patke et al., 2024; Yu et al., 2025), which presents an opportunity for the parameter remapping engine to prioritize remapping parameters of inactive models. However, this approach requires careful design, as overly aggressive offloading may lead to starvation – especially given the unpredictable nature of request arrival patterns in online serving environments (Stojkovic et al., 2024b). An inefficient design is likely to prevent certain models from responding promptly; for example, a suboptimal model selection can lead to a 22.0% increase in serving tail latency (§ 7.6).
@@
How many layers to remap? Determining the optimal remapping percentage is challenging due to the dynamic nature of request arrival rates in real-world online inference serving (Stojkovic et al., 2024a; Patel et al., 2024; Stojkovic et al., 2024b). The fluctuations in request arrival rates can affect the GPU computation time, which in turn affects the percentage of parameters that can be transferred without performance degradation. We illustrate the need for dynamic adjustment of the remapping percentage using an example of serving OPT-13b with varying batch sizes on the ShareGPT dataset (Team, 2025). Figure 4 shows how the GPU computation time for generating one token in the decode phase varies with batch size, while parameter load time – determined solely by the remapping percentage – remains constant. The intersection points of these curves represent configurations where parameter loading is fully overlapped with computation, maximizing efficiency. As shown, the optimal remapping percentage varies with changes in the runtime batch size, reflecting shifts in system load. Using a static remapping percentage may result in suboptimal resource use: underutilizing GPU memory at off-peak load and causing compute stalls at peak load. For example, a 30% remapping percentage tuned for average batch size in ShareGPT results in CPU-GPU bandwidth becoming a bottleneck during off-peak workloads. Therefore, the remapping percentage must be dynamically adapted at runtime.
@@
Which layers to target? Once the remapping percentage is set, the next design challenge is selecting specific layers to remap. This remains non-trivial, as Oneiros must decide how to allocate limited GPU memory across all layers – deciding which layers remain permanently in GPU memory and which are transferred on the fly. For an efficient pipeline, a layer selection strategy is needed to guide the layer-sharing mechanism in Oneiros. A suboptimal layer selection can lead to reduced serving throughput (by 12.7%, § 7.5).
4. Oneiros Overview
@@
We introduce Oneiros, a Dynamic Remapping Engine for multi-tenant LLM inference serving systems that elastically adjusts the KV cache size by remapping parameter memory as KV cache memory. Figure 6 shows the Oneiros architecture, and depicts how Oneiros can be integrated into existing LLM inference serving systems.
4.1. Architecture of Oneiros
@@
Oneiros is designed to have three key components, marked in green color in Figure 6.
@@
Async Transfer Engine: The Async Transfer Engine manages parameter loading and synchronizes CPU-GPU data transfers with GPU computation. When the Remapping Controller initiates parameter remapping, a portion of GPU memory originally allocated for model parameters is repurposed for KV cache usage. If the reclaimed parameters are later required for computation, the Async Transfer Engine fetches them back from CPU memory to GPU memory. By leveraging the deterministic execution order of layers within each decode iteration, the Async Transfer Engine overlaps CPU-GPU data transfers with ongoing GPU computation.
Figure 6. Overview of Oneiros.
@@ 4.2. Workflow
Figure 6 depicts the workflow of the interaction between various components to achieve the goal of dynamic parameter remapping in Oneiros. The Scheduler selects which models to serve, referred to as active models. Once active models are chosen, the Memory Allocator updates the parameter and KV cache memory usage accordingly. Using both model information and KV cache utilization data, the Remapping Controller determines whether parameter memory needs to be remapped to expand KV cache capacity. If remapping is required, the Remapping Controller updates memory usage records in the Memory Allocator and triggers the Async Transfer Engine to initiate data transfer. The GPU then executes the LLM inference tasks. With support from Oneiros, a portion of parameter memory in GPUs can be dynamically reclaimed at runtime, allowing the system to accommodate larger batches or longer sequences.
@@ 5. Remapping Controller
Oneiros enables dynamic, runtime management of GPU memory through its Remapping Controller. Figure 7 illustrates an example execution workflow of the Remapping Controller across five phases, each representing a snapshot of model activity within an LLM serving system. Depending on runtime conditions, the Remapping Controller decides either to initiate remapping, continue remapping additional layers from the active or inactive models, or halt remapping altogether (reclaimed KV cache memory restored for parameter storage). In this section, we describe how Remapping Controller answers the key questions, including when to remap (§ 5.1), which models to remap (§ 5.2), how many (§ 5.3), and which layers to remap (§ 5.4). Putting it all together, we then present an algorithm that can seamlessly integrate Oneiros into existing LLM inference serving frameworks, enabling dynamic remapping of model parameter memory to automatically and elastically expand KV cache capacity at runtime (§ 5.5).
@@ 5.1. When to Remap?
The Remapping Controller initiates parameter remapping when the KV cache capacity is exhausted. While unidirectional parameter remapping incurs lower overhead than bidirectional KV cache swapping, Oneiros activates it only when necessary. As such, accurately determining both the initiation and termination points of remapping is critical. Given the dynamic nature of real-world production workloads (Cortez et al., 2017; Hadary et al., 2020; Shahrad et al., 2020; Stojkovic et al., 2024a; Patel et al., 2024; Stojkovic et al., 2024b), the Remapping Controller enables remapping during peak request periods to accommodate larger batch sizes. During non-peak periods, when KV cache space is sufficient, remapping is disabled to avoid the overhead of reloading previously evicted parameters into GPU memory. The reclaimed KV cache memory is then restored for parameter usage.
@@ 5.2. Which Models to Remap?
GPU sharing background: Given the memory capacity of emerging GPUs, enabling GPU sharing is essential to fully utilize their processing capabilities. Prior works have examined both temporal (Bai et al., 2020; Gujarati et al., 2020; Xiao et al., 2020; Fu et al., 2024; Patke et al., 2024) (example in Figure 5) and spatial (Dhakal et al., 2020; Romero et al., 2021; Lee et al., 2024b; Strati et al., 2024; Choi et al., 2022; Xu et al., 2022; Han et al., 2024a; Wang et al., 2024a) (example in Figure 5) sharing strategies for efficient machine learning inference serving on GPUs. Temporal sharing allocates the entire GPU to one model at a time, allowing for larger batches per model. This approach is especially useful in multi-agent workflows – where the output of one model becomes the input to another (Talebirad and Nadiri, 2023; Chan et al., 2023; Li et al., 2024b; Han et al., 2024b; Kim et al., 2025) – and in cloud production environments, where models often experience idle periods with no incoming requests (Yu et al., 2025). In contrast, spatial sharing divides GPU resources among multiple models simultaneously, allowing for faster response times at the expense of smaller batch sizes per model (Lee et al., 2024b; Duan et al., 2024). Each GPU resource sharing strategy is tailored to specific use cases and can be selectively employed by multi-tenant LLM inference schedulers. The Remapping Controller in Oneiros is designed to be compatible with any scheduling policy, although the design of the scheduler itself is beyond the scope of this study.
Model selection strategy: We now present the model selection strategy to determine which models to remap parameter layers from. The Remapping Controller prioritizes remapping parameters from inactive models with the lowest priority, as determined by the scheduling policy, rather than from active ones. Under temporal GPU resource sharing, models scheduled for future execution are considered inactive. In contrast, under spatial sharing, all models are treated as active, requiring each to remap parameters. The number of layers to remap and their selection are determined using the strategies described in § 5.3 and § 5.4. If the scheduler does not specify model priorities or all models are active, the Remapping Controller defaults to a round-robin policy, treating all models equally, thereby ensuring fairness across models. In this case, it applies a Most Recently Used (MRU) strategy, remapping parameters from the most recently activated model to defer parameter transfer costs as far into the future as possible. To avoid any starvation, Oneiros also caps the maximum number of layers remapped based on a threshold derived empirically through runtime profiling (more in § 5.3), and no model is ever remapped in its entirety. This strategy is especially effective when future request patterns are unpredictable or when model priorities are unspecified.
@@ 5.3. How Many Layers to Remap?
To hide parameter transfer latency by overlapping it with GPU computation, the transfer from CPU to GPU must finish before the computation completes; otherwise, it becomes a performance bottleneck. Since both model parameter sizes and CPU-GPU bandwidth are known ahead of runtime, the transfer time per layer can be profiled offline and treated as prior knowledge by the online inference system. As a result, the Remapping Controller only needs to monitor GPU computation time, denoted as $T_{Compute}$, at runtime to determine whether additional parameters can be remapped. For an active model, $T_{Compute}$ refers to the time required to generate the current token during the decode phase. For an inactive model, it represents the duration of the prefill phase, which is more compute-intensive and typically longer, as no tokens have been generated yet (Patel et al., 2024; Agrawal et al., 2024). Let $T_{T}$ denote the time required to transfer a single layer of parameters. To prevent parameter loading from becoming a new bottleneck, the number of remapped layers $N$ must satisfy the constraint $T_{T}\cdot N\leq T_{Compute}$.
The longer $T_{Compute}$ in the prefill phase allows more layers in inactive models to be remapped compared to when the model is active. Similar to the decode phase, the prefill phase executes one layer at a time rather than all at once. If all the parameters are not present in GPU memory because of the parameter remapping triggered by Oneiros, the Remapping Controller transfers the parameters for the later layers from CPU memory back into the GPU memory while initial layers are being executed. Thus, the parameters are ready in GPU memory when needed, as the prefill phase makes progress layer-by-layer. The prefill phase is compute-bound, and takes additional compute time that provides an opportunity for Oneiros to hide the latency of loading remapped parameters back to GPU memory.
@@ 5.4. Which Layers to Target?
Layer selection strategy: Understanding the layer selection strategy requires recognizing the circular execution pattern of layers in LLM inference, as shown in Figure 7. The auto-regressive nature of LLM inference makes each of the subsequent tokens execute the same layers again. Thus, LLM inference can be viewed as a circular execution of layers, where the execution of the last layer for the current token is succeeded by the execution of the first layer for the next token. The Remapping Controller uses a uniform-interval layer selection strategy, in which evenly spaced layers are chosen to share a common GPU memory region allocated for their parameters. These layers alternate access to the shared region – at any given moment, some layer parameters reside in GPU memory, while others remain in CPU memory. When a layer not currently in GPU memory is needed, its parameters are loaded into the shared region, replacing those of the previously loaded layer.
@@
Example: To provide a more intuitive explanation, we use the example of Model-C with 8 layers (Phase 2 of Figure 7). The Remapping Controller employs a uniform-interval layer selection strategy. For instance, when remapping the parameters of a single layer, Layers 1 and 5 are selected to share a common GPU memory region (the other 6 layers remain in GPU memory permanently), alternating between CPU and GPU memory. The uniform-interval layer selection strategy is designed to balance serving latency across successive token generations. In contrast, ignoring the circular nature of LLM inference – by treating each token generation as an independent execution – may lead to suboptimal choices, such as selecting Layer 1 and Layer 8. Although this configuration allows sufficient time to load the parameters of Layer 8 after executing Layer 1, it does not provide enough time to reload the parameters of Layer 1 before it is needed again. This demonstrates the benefit of uniform-interval selection in circular execution, which ensures that each remapped layer can be restored to GPU memory in time for reuse.
@@
Theoretical proof: Consider an n-layer model in which the Remapping Controller remaps the parameter memory of $\alpha$ layers for KV cache. Consequently, all $n$ layers must be scheduled within the GPU memory footprint of only $n-\alpha$ layers. Suppose $m$ layers are transferred between CPU and GPU during each token generation. We note $L_{1},L_{2},\ldots,L_{m}$ as the $m$ layers (remapped layers in Figure 7) that are transferred from CPU to GPU. We define $T_{T}$ as the time taken to load parameters of a single layer from CPU to GPU memory, and $T_{c}$ as the time spent on computation for a single layer on GPU. Our layer selection policy aims to satisfy two key objectives: (1) ensure that parameter transfer latency does not become a performance bottleneck; and (2) minimize CPU-GPU data transfer traffic.
@@
Consider that there are $k_{1}$ layers between $L_{1}$ and $L_{2}$ [^4] In this paper, we assume all layers are identical. For emerging models with heterogeneity across layers (Zhang et al., 2025), these formulas can be easily adapted, e.g., $T_{T}$ can be written as $T_{T,i}$, where $i$ denotes each specific layer, and solved using integer linear programming., and $k_{m}$ layers between $L_{m}$ and $L_{1}$ (for the next token), we have: